# Properties of kaon-condensed phase in hyperon-mixed matter with three-baryon forces

### v1

@@ VII.1 Composition in the ($Y$+$K$) phase
In case KC sets in at a lower density than the $\Xi^{-}$ hyperons (for $L$ = 60 MeV with $\Sigma_{Kn}$ = 300 MeV, or for $\Sigma_{Kn}$ = 400 MeV), mixing of the $\Xi^{-}$ hyperons is pushed up to high densities, or even does not occur over the relevant densities, as seen from Figs. 89, and Table 3. (Note that for the case of $L$ = 70 MeV with $\Sigma_{Kn}$ = 400 MeV, the $\Xi^{-}$ hyperons precede KC, but their fraction soon vanishes, until they reappear at high density, $\rho_{\rm B}$ = 0.79 fm-3.) On the contrary, for $L$ = 65 MeV and 70 MeV with $\Sigma_{Kn}$ = 300 MeV, $\Xi^{-}$ hyperons appear at a lower density than KC, and the onset of KC is pushed up to high densities. Due to the repulsive $K$-$\Xi^{-}$ vector interaction term in $X_{0}$, the form of which is specified by chiral symmetry, KC and $\Xi^{-}$ hyperons compete against each other. This competitive relationship can also be seen from the interaction part of the $K^{-}$ number density $\rho_{K^{-}}$ [Eq. (70)]. Thus the existence of KC strongly affects onset of the $\Xi^{-}$ hyperons and subsequent particle composition at high densities. In Table 3, the onset densities at which hyperon ($\Lambda$, $\Xi^{-}$)-mixing starts and those of KC are summarized. It is to be noted that the $\Sigma^{-}$ hyperons are not mixed over the relevant densities due to the strong repulsion of the $V_{\Sigma^{-}}^{N}$ in our model.
@@ VII.3 Energy per baryon and pressure
It should be mentioned here that there is a causal limit, beyond which the sound velocity exceeds the speed of light, in our model for each case of $L$, specifically beyond $\rho_{\rm B}\sim$ 1.03 fm-3. As indicated in Fig. 14, the causality condition is fulfilled for a larger value of $L$ = 70 MeV until the mass reaches the maximum mass, while it is violated for $L$ = 60 MeV and 65 MeV before the mass reaches the maximum mass. The many-baryon repulsion more than two-body force eventually leads to violation of the causality condition for sufficiently high baryon densities. In order to improve the EOS so as to meet the causality condition over the whole densities in compact stars, we adopted the Lorentz scalar form of the many-baryon repulsive force and tried to obtain the EOS with the ($Y$+$K$) phase which is sufficient for the fulfillment of causality condition muto2024 . It is a common feature that the sound velocity $v_{s}$ in hadronic matter becomes large as density increases and that it approaches the light velocity $c$ due to the many-baryon repulsive forces. The result should be compared with the case of hadron-quark crossover baym2019 ; kojo2021 ; fujimoto2022 .
@@ VIII.1 Gravitational mass to radius and central density relations
Here we discuss the effects of KC on the static properties of compact stars. In Fig. 14, the gravitational mass $M$ radius $R$ relations are shown for $L$ = (60, 65, 70) MeV. In Fig. 14, the gravitational mass $M$ is shown as functions of baryon number density at the center of the star, $\rho_{\rm B,center}$. They are obtained by solving the Tolman-Oppenheimer-Volkoff equation with the EOS including the ($Y$+$K$) phase. For low density region $\rho_{\rm B}<$ 0.10 fm-3 below the density of uniform matter, we utilize the EOS of Ref. BPS1971 and combine with the EOS obtained in our model for $\rho_{\rm B}\geq$ 0.10 fm-3.
@@
The branches including KC in the core are denoted as the bold solid lines (thin solid lines) for $\Sigma_{Kn}$ = 300 MeV (400 MeV). For comparison, the branch including pure hyperon-mixed matter, where KC is switched off by setting $\theta=0$, is shown by the green dashed line for each case of $L$. The filled triangle [$\blacktriangle$] stands for the branch point where the $\Lambda$ hyperons appear from nuclear matter in the center of the star. The branch point at which KC appears in the center of the star is indicated by the filled circle [$\bullet$] (open circle [$\circ$]) in the case of $\Sigma_{Kn}$ = 300 MeV (400 MeV). In Fig. 14, the location of the maximum mass point for each branch including the ($Y$+$K$) phase is indicated by the open square [$\square$]. The cross point [$\times$] corresponds to the causal limit beyond which the sound velocity exceeds the speed of light. Figure 14 shows that the maximum mass and its radius shift to larger values for larger $L$. Also the radius of neutron stars for a given mass in the stable branch increases with $L$. These properties reflect that the EOS becomes stiffer at high densities for larger $L$, due to the fact that the repulsive contribution from the “two-body” baryon interaction gets large for large $L$ (VII.3). One can also see the $L$-dependence of the stiffness of the EOS from Fig. 14. At a given $M$, the density at the center of the star, $\rho_{\rm B,center}$, is smaller for larger $L$.
@@
The curves of $M$-$R$ relations based on our EOS in the case of $L$ =(65, 70) MeV pass through the above constrained regions. In particular, the maximum masses with the ($Y$+$K$) phase in the core are consistent with recent observations of massive neutron stars in both cases of $\Sigma_{Kn}$ = 300 MeV and 400 MeV for $L$ =(65, 70) MeV. However, the masses within the causal limit for $\Sigma_{Kn}$ = 400 MeV and $L$ = 60 MeV do not reach the range allowable from the observations of most massive neutron stars. In our model, the larger values of the slope $L>$ 60 MeV are preferred in order to obtain observed massive neutron stars. The radius of PSR J0740+6620 has recently been improved by the use of the updated NICER data and the $X$-ray Multi-Mirror (XMM-Newton) observation salmi2024 ; dittmann2024 . Our result for $L$ =(65, 70) MeV is still consistent with this new constraint.
@@
For neutron stars with $M\lesssim$1.4 $M_{\odot}$, the central density does not reach $\rho_{\rm B}^{c}(\Lambda)\sim 0.4$ fm-3 for $L$=(65, 70) MeV, and the ground state in the core consists of only $n$, $p$, and leptons ($e^{-}$, $\mu^{-}$).
In relation to a neutron star cooling, we show, with the asterisk in Fig. 14, the threshold density at the center of the star for the direct Urca process ($np$-DU), $n\rightarrow p+l+\bar{\nu}_{l}$, $p+l\rightarrow n+\nu_{l}$ ($l=e^{-},\mu^{-}$lattimer1991 , and corresponding gravitational mass of the whole star. For a star with canonical mass 1.4 $M_{\odot}$, the central density $\rho_{\rm B,center}$ slightly exceeds the threshold density for the $np$-DU process for $L$ = (60--70) MeV.
For a heavy neutron star with $\sim$ 2.0 $M_{\odot}$, one can expect the ($Y$+$K$) phase in the core of the star. In the presence of KC, extra rapid cooling processes by neutrino and antineutrino emissions, i. e. the kaon-induced Urca (KU) process for nucleon ($N=p,n$), $N+\langle K^{-}\rangle\rightarrow N+l+\bar{\nu}_{l}$, $N+l\rightarrow N+\langle K^{-}\rangle+\nu_{l}$, can occur with $\langle K^{-}\rangle$ being the classical $K^{-}$ field t88 ; bkpp1988 . In the ($Y$+$K$) phase, the kaon-induced Urca process for hyperons (abbreviated to $Y$-KU) may also be suggested, while the dominant processes may be the hyperon ($\Lambda$) direct Urca process plp1992 from the region surrounding the ($Y$+$K$) core. These weak processes are important to investigate the possible existence of the ($Y$+$K$) phase.
@@ IX Summary and concluding remarks
We have investigated the coexistent phase of kaon condensates and hyperons [($Y$+$K$) phase] by the use of the interaction model based on the effective chiral Lagrangian (ChL) for $K$-$B$ and $K$-$K$ interactions combined with the minimal relativistic mean-field theory (MRMF) for two-body baryon interaction, taking into account the universal three-baryon repulsion (UTBR) and the phenomenological three-nucleon attraction (TNA). Interplay between KC and hyperons and resulting onset mechanisms of KC in hyperon-mixed matter and the EOS with the ($Y$+$K$) phase have been clarified within the (ChL+MRMF+UTBR+TNA) model. While the $s$-wave $K$-$B$ scalar attraction simulated by the $K$-$B$ sigma terms favors coexistence of KC and hyperons, they tend to compete with each other by the $s$-wave $K$-$B$ vector repulsion as well as diminishing the $K$-$N$ ($N=p,n$) vector attraction as a result of mixing of hyperons (especially $\Lambda$ hyperons). In addition, the former attraction gets moderate at high densities because of suppression of the total baryon scalar density as the relativistic effect. Further, the $K$-$B$ scalar attraction is suppressed more along with the growth of KC, which we call self-suppression mechanism. As a result, softening of the EOS due to KC becomes moderate in the presence of hyperons at high densities.
@@
The EOS and the resulting mass and radius of compact stars within hadronic picture accompanying the ($Y$+$K$) phase are consistent with recent observations of massive neutron stars. In our interaction model (ChL+MRMF+UTBR+TNA), the higher value of the slope $L$ ($\gtrsim$ 65 MeV) is preferred, since, for $L<$ 60MeV, considerable softening of the EOS due to the appearance of hyperons and/or KC at high densities ($\rho_{\rm B}\gtrsim 2.5\rho_{0}$) cannot be compensated with the stiffening effect due to the Two-BR.
We have assumed that the UTBR works between different kinds of baryons on an equal footing on the ground that such repulsion at high densities originates from confinement of quarks in the baryons and is independent of spin-flavor of baryons. Validity of the universality of the UTBR should be examined by comparing with other results in chiral perturbation theory applied to baryons petschauer2020 , quark models including the quark Pauli effects oka2012 ; nakamoto2016 and lattice QCD results inoue2019 .
It has been suggested from the transverse and elliptic flow data in nuclear collisions that some extra softening may be needed for the EOS over baryon densities (2--4)$\rho_{0}$ danielewicz2002 . One of the feasible solutions is appearance of pion condensation (PC) driven by the $p$-wave $\pi NN$ attractive interaction. The onset density of charged pion ($\pi^{c}$) condensation with realistic medium effects is estimated to be $\rho_{\rm B}\sim 2\rho_{0}$ tatsumi1982 , so that the combined condensation of $\pi^{c}$ and kaons ($\pi$-$K$ condensation) at high densities should be elaborated within the relativistic framework. In the combined $\pi$-$K$ condensation, not only PC but also KC have a spatial momentum through the $p$-wave $\pi$-$K$ interactions. In the ground state of the $\pi$-$K$ condensed phase in neutron-star matter without hyperon-mixing, the energy eigenstates are given by quasi-baryonic states with superposition of neutron and proton states under the $p$-wave $\pi^{c}$ condensates. The ground state is occupied solely by the lower energy eigenstates of the quasi-baryons, forming the one-Fermi sea, which may resolve the assumption of the universal strengths between different species of baryons for the UTBR.
@@
As stated in Sec. VII.3, the EOS in our model faces with a causal limit at $\rho_{\rm B}\sim$1.03 fm-3. In the recent work, we have improved the UTBR by assuming a Lorentz-scalar form so as to safely meet the causality condition over the relevant densities realized in compact stars muto2024 . Another scenario for preventing the hadronic EOS from causality violation is a possible connection of hadronic matter including the ($Y$+$K$) phase to quark matter at high densities. In particular, there are extensive studies in which hadronic matter was connected to quark matter smoothly by crossover transition to obtain massive neutron stars compatible with observations mht2013 ; baym2018 ; baym2019 ; kojo2021 ; fujimoto2022 . In our case, the baryon density at causal limit ($\sim$ 1.03 fm-3) corresponds to the density where repulsive core of baryons being set with 0.5 fm-3 is attached together, implying the beginning of quark percolation. The connection of hadronic phase including the ($Y$+$K$) phase with quark degrees of freedom at high densities is needed in future works. Role of meson condensation in both hadron phase and quark phase is also an important subject.
In order to clarify the existence of the ($Y$+$K$) phase, novel features associated with the ($Y$+$K$) phase and their implications for astrophysical phenomena should be explored. For example, KC might reveal unique features as superconductivity through responses to rotation, strong magnetic field, and radial/nonradial oscillations of compact stars as a result of strong coupling to baryonic system. With regard to thermal evolution of neutron stars, several neutron stars have anomalously low temperature that suggests extraordinary rapid cooling processes tsuruta1998 ; dohi2019 ; dohi2022 ; marino2024 . It has been shown that the relevant weak reactions and $K$-$B$ dynamics associated with KC can be described in a unified manner on the basis of chiral symmetry t88 . The emissivities for these reactions should be obtained in conjunction with the quantities such as composition of matter for the ($Y$+$K$) phase. Cooling curves with the use of the emissivity in the ($Y$+$K$) phase and comparison with observation of surface temperatures should be elucidated in detail bdmn2024 .

### v2

@@ VII.1 Composition in the ($Y$+$K$) phase
In case KC sets in at a lower density than the $\Xi^{-}$ hyperons (for $L$ = 60 MeV with $\Sigma_{Kn}$ = 300 MeV, or for $\Sigma_{Kn}$ = 400 MeV), mixing of the $\Xi^{-}$ hyperons is pushed up to high densities, or even does not occur over the relevant densities, as seen from Figs. 89, and Table 3. (Note that for the case of $L$ = 70 MeV with $\Sigma_{Kn}$ = 400 MeV, the $\Xi^{-}$ hyperons precede KC, but their fraction soon vanishes, until they reappear at high density, $\rho_{\rm B}$ = 0.79 fm-3.) On the contrary, for $L$ = 65 MeV and 70 MeV with $\Sigma_{Kn}$ = 300 MeV, $\Xi^{-}$ hyperons appear at a lower density than KC, and the onset of KC is pushed up to high densities. Due to the repulsive $K$-$\Xi^{-}$ vector interaction term in $X_{0}$, the form of which is specified by chiral symmetry, KC and $\Xi^{-}$ hyperons compete against each other. This competitive relationship can also be seen from the interaction part of the $K^{-}$ number density $\rho_{K^{-}}$ [Eq. (70)]. Thus the existence of KC strongly affects onset of the $\Xi^{-}$ hyperons and subsequent particle composition at high densities. In Table 3, the onset densities at which hyperon ($\Lambda$, $\Xi^{-}$)-mixing starts and those of KC are summarized. It is to be noted that the $\Sigma^{-}$ hyperons are not mixed over the relevant densities due to the strong repulsion of the $V_{\Sigma^{-}}^{N}$ in the model.
@@ VII.3 Energy per baryon and pressure
In contrast, the repulsive energy stemming from the two-body $B$-$B$ interaction becomes always dominant over the repulsive contribution from the NLSI terms at high densities regardless of the choice of $L$ in the (MRMF+NLSI) model. This is because the many-baryon repulsion tends to diminish the meson mean-fields at high densities, leading to an energetically stable state. [See Fig. 17 in Appendix X.4 in the case of $L$ = 65 MeV for the SNM, and also Fig. 6 in Ref. mmt2022 for hyperon-mixed matter.]
@@
It should be mentioned here that there is a causal limit, beyond which the sound velocity exceeds the speed of light, in the present model for each case of $L$, specifically beyond $\rho_{\rm B}\approx$ 1.03 fm-3. As indicated in Fig. 14, the causality condition is fulfilled for a larger value of $L$ = 70 MeV until the mass reaches the maximum mass, while it is violated for $L$ = 60 MeV and 65 MeV before the mass reaches the maximum mass. [See also Table 3 for the central density of the maximum mass star, $\rho_{\rm B,center}(M_{\rm max})$.] The many-baryon repulsion more than two-body force eventually leads to violation of the causality condition for sufficiently high baryon densities. In order to improve the EOS so as to meet the causality condition over the whole densities in compact stars, one adopted the Lorentz scalar form of the many-baryon repulsive force and tried to obtain the EOS with the ($Y$+$K$) phase which is sufficient for the fulfillment of causality condition muto2024 . It is a common feature that the sound velocity $v_{s}$ in hadronic matter becomes large as density increases and that it approaches the light velocity $c$ due to the many-baryon repulsive forces. The result should be compared with the case of hadron-quark crossover baym2019 ; kojo2021 ; fujimoto2022 .
@@ VIII.1 Gravitational mass to radius and central density relations
Here the effects of KC on the static properties of compact stars are discussed. In Fig. 14, the gravitational mass $M$ radius $R$ relations are shown for $L$ = (60, 65, 70) MeV. In Fig. 14, the gravitational mass $M$ is shown as functions of baryon number density at the center of the star, $\rho_{\rm B,center}$. They are obtained by solving the Tolman-Oppenheimer-Volkoff equation with the EOS including the ($Y$+$K$) phase. For low density region $\rho_{\rm B}<$ 0.10 fm-3 below the density of uniform matter, one utilizes the EOS of Ref. BPS1971 and combine with the EOS obtained in the model for $\rho_{\rm B}\geq$ 0.10 fm-3.
@@
The branches including KC in the core are denoted as the bold solid lines (thin solid lines) for $\Sigma_{Kn}$ = 300 MeV (400 MeV). For comparison, the branch including pure hyperon-mixed matter, where KC is switched off by setting $\theta=0$, is shown by the green dashed line for each case of $L$. The filled triangle [$\blacktriangle$] stands for the branch point where the $\Lambda$ hyperons appear from nuclear matter in the center of the star. The branch point at which KC appears in the center of the star is indicated by the filled circle [$\bullet$] (open circle [$\circ$]) in the case of $\Sigma_{Kn}$ = 300 MeV (400 MeV). In Fig. 14, the location of the maximum mass point for each branch including the ($Y$+$K$) phase is indicated by the open square [$\square$]. The cross point [$\times$] corresponds to the causal limit beyond which the sound velocity exceeds the speed of light. Figure 14 shows that the maximum mass and its radius shift to larger values for larger $L$. Also the radius of neutron stars for a given mass in the stable branch increases with $L$. These properties reflect that the EOS becomes stiffer at high densities for larger $L$, due to the fact that the repulsive contribution from the “two-body” baryon interaction gets large for large $L$ (VII.3). One can also see the $L$-dependence of the stiffness of the EOS from Fig. 14. At a given $M$, the density at the center of the star, $\rho_{\rm B,center}$, is smaller for larger $L$. For reference, the central density of the maximum mass star, $\rho_{\rm B,center}(M_{\rm max})$, is listed in Table 3.
@@
The curves of $M$-$R$ relations based on the EOS in the case of $L$ =(65, 70) MeV pass through the above constrained regions. In particular, the maximum masses with the ($Y$+$K$) phase in the core are consistent with recent observations of massive neutron stars in both cases of $\Sigma_{Kn}$ = 300 MeV and 400 MeV for $L$ =(65, 70) MeV. However, the masses within the causal limit for $\Sigma_{Kn}$ = 400 MeV and $L$ = 60 MeV do not reach the range allowable from the observations of most massive neutron stars. In this model, the larger values of the slope $L>$ 60 MeV are preferred in order to obtain observed massive neutron stars. The radius of PSR J0740+6620 has recently been improved by the use of the updated NICER data and the $X$-ray Multi-Mirror (XMM-Newton) observation salmi2024 ; dittmann2024 . The result for $L$ =(65, 70) MeV is still consistent with this new constraint.
@@
For neutron stars with $M\lesssim$1.4 $M_{\odot}$, the central density does not reach $\rho_{\rm B}^{c}(\Lambda)\approx 0.4$ fm-3 for $L$=(65, 70) MeV, and the ground state in the core consists of only $n$, $p$, and leptons ($e^{-}$, $\mu^{-}$).
In relation to a neutron star cooling, one shows, with the asterisk in Fig. 14, the threshold density at the center of the star for the direct Urca process ($np$-DU), $n\rightarrow p+l+\bar{\nu}_{l}$, $p+l\rightarrow n+\nu_{l}$ ($l=e^{-},\mu^{-}$lattimer1991 , and corresponding gravitational mass of the whole star. For a star with canonical mass 1.4 $M_{\odot}$, the central density $\rho_{\rm B,center}$ slightly exceeds the threshold density for the $np$-DU process for $L$ = (60--70) MeV.
For a heavy neutron star with $\approx$ 2.0 $M_{\odot}$, one can expect the ($Y$+$K$) phase in the core of the star. In the presence of KC, extra rapid cooling processes by neutrino and antineutrino emissions, i. e. the kaon-induced Urca (KU) process for nucleon ($N=p,n$), $N+\langle K^{-}\rangle\rightarrow N+l+\bar{\nu}_{l}$, $N+l\rightarrow N+\langle K^{-}\rangle+\nu_{l}$, can occur with $\langle K^{-}\rangle$ being the classical $K^{-}$ field t88 ; bkpp1988 . In the ($Y$+$K$) phase, the kaon-induced Urca process for hyperons (abbreviated to $Y$-KU) may also be suggested, while the dominant processes may be the hyperon ($\Lambda$) direct Urca process plp1992 from the region surrounding the ($Y$+$K$) core. These weak processes are important to investigate the possible existence of the ($Y$+$K$) phase.
@@ IX Summary and concluding remarks
One has investigated the coexistent phase of kaon condensates and hyperons [($Y$+$K$) phase] by the use of the interaction model based on the effective chiral Lagrangian (ChL) for $K$-$B$ and $K$-$K$ interactions combined with the minimal relativistic mean-field theory (MRMF) for two-body baryon interaction, taking into account the universal three-baryon repulsion (UTBR) and the phenomenological three-nucleon attraction (TNA). Interplay between KC and hyperons and resulting onset mechanisms of KC in hyperon-mixed matter and the EOS with the ($Y$+$K$) phase have been clarified within the (ChL+MRMF+UTBR+TNA) model. While the $s$-wave $K$-$B$ scalar attraction simulated by the $K$-$B$ sigma terms favors coexistence of KC and hyperons, they tend to compete with each other by the $s$-wave $K$-$B$ vector repulsion as well as diminishing the $K$-$N$ ($N=p,n$) vector attraction as a result of mixing of hyperons (especially $\Lambda$ hyperons). In addition, the former attraction gets moderate at high densities because of suppression of the total baryon scalar density as the relativistic effect. Further, the $K$-$B$ scalar attraction is suppressed more along with the growth of KC, which is called as self-suppression mechanism. As a result, softening of the EOS due to KC becomes moderate in the presence of hyperons at high densities.
@@
The EOS and the resulting mass and radius of compact stars within hadronic picture accompanying the ($Y$+$K$) phase are consistent with recent observations of massive neutron stars. In the interaction model (ChL+MRMF+UTBR+TNA), the higher value of the slope $L$ ($\gtrsim$ 65 MeV) is preferred, since, for $L<$ 60MeV, considerable softening of the EOS due to the appearance of hyperons and/or KC at high densities ($\rho_{\rm B}\gtrsim 2.5\rho_{0}$) cannot be compensated with the stiffening effect due to the Two-BR.
One has assumed that the UTBR works between different kinds of baryons on an equal footing on the ground that such repulsion at high densities originates from confinement of quarks in the baryons and is independent of spin-flavor of baryons. Validity of the universality of the UTBR should be examined by comparing with other results in chiral perturbation theory applied to baryons petschauer2020 , quark models including the quark Pauli effects oka2012 ; nakamoto2016 and lattice QCD results inoue2019 .
It has been suggested from the transverse and elliptic flow data in nuclear collisions that some extra softening may be needed for the EOS over baryon densities (2--4)$\rho_{0}$ danielewicz2002 . One of the feasible solutions is appearance of pion condensation (PC) driven by the $p$-wave $\pi NN$ attractive interaction. The onset density of charged pion ($\pi^{c}$) condensation with realistic medium effects is estimated to be $\rho_{\rm B}\approx 2\rho_{0}$ tatsumi1982 , so that the combined condensation of $\pi^{c}$ and kaons ($\pi$-$K$ condensation) at high densities should be elaborated within the relativistic framework. In the combined $\pi$-$K$ condensation, not only PC but also KC have a spatial momentum through the $p$-wave $\pi$-$K$ interactions. In the ground state of the $\pi$-$K$ condensed phase in neutron-star matter without hyperon-mixing, the energy eigenstates are given by quasi-baryonic states with superposition of neutron and proton states under the $p$-wave $\pi^{c}$ condensates. The ground state is occupied solely by the lower energy eigenstates of the quasi-baryons, forming the one-Fermi sea, which may resolve the assumption of the universal strengths between different species of baryons for the UTBR.
@@
As stated in Sec. VII.3, the EOS in the present model faces with a causal limit at $\rho_{\rm B}\approx$1.03 fm-3. In the recent work, one has improved the UTBR by assuming a Lorentz-scalar form so as to safely meet the causality condition over the relevant densities realized in compact stars muto2024 . Another scenario for preventing the hadronic EOS from causality violation is a possible connection of hadronic matter including the ($Y$+$K$) phase to quark matter at high densities. In particular, there are extensive studies in which hadronic matter was connected to quark matter smoothly by crossover transition to obtain massive neutron stars compatible with observations mht2013 ; baym2018 ; baym2019 ; kojo2021 ; fujimoto2022 . In case of the model in this paper, the baryon density at causal limit ($\approx$ 1.03 fm-3) corresponds to the density where repulsive core radius of baryons being set to 0.5 fm is attached together, implying the beginning of quark percolation. The connection of hadronic phase including the ($Y$+$K$) phase with quark degrees of freedom at high densities is needed in future works. Role of meson condensation in both hadron phase and quark phase is also an important subject.
In order to clarify the existence of the ($Y$+$K$) phase, novel features associated with the ($Y$+$K$) phase and their implications for astrophysical phenomena should be explored. For example, the effect of strong magnetic field on the onset density of KC and the EOS in neutron-star matter has been studied with the quark-meson coupling model ys2008 . It has been shown that strong magnetic field as large as $10^{18}$ G may alter the onset density of KC, matter composition, and the EOS significantly. For another example, effects of the symmetry energy on KC were studied for protoneutron stars with fixed entropy and trapped neutrinos with the modified quark-meson coupling model. It has been suggested that the melting of KC and a second formation of KC may be signalized by neutrinos during the cooling process panda2014 . Further, KC might reveal unique features as superconductivity through responses to rotation, strong magnetic field, and radial/nonradial oscillations of compact stars as a result of strong coupling to baryonic system.
With regard to thermal evolution of neutron stars, several neutron stars have anomalously low temperature that suggests extraordinary rapid cooling processes tsuruta1998 ; dohi2019 ; dohi2022 ; marino2024 . It has been shown that the relevant weak reactions and $K$-$B$ dynamics associated with KC can be described in a unified manner on the basis of chiral symmetry t88 . The emissivities for these reactions should be obtained in conjunction with the quantities such as composition of matter for the ($Y$+$K$) phase. Cooling curves with the use of the emissivity in the ($Y$+$K$) phase and comparison with observation of surface temperatures should be elucidated in detail bdmn2024 .